Why do parallel circuits have maximum energy compared to series circuits?
Answer #1

Suppose two resistors, each of resistance R are connected in parallel to battery of voltage V. Then the voltage drop across each resistor is V. The power delivered or energy delivered per unit time across each resistor is

The total power delivered by the battery to both the resistors is

You can also derive the result by considering the net resistance of the parallel circuit which is 1/(1/R+1/R) = R/2

Hence the total power delivered by the battery to both the resistors in parallel is

Supppose these two resistors, each of resistance R is connected in series to the battery of voltage V. The total resistance is R+R = 2R

The total power delivered by the battery to both the resistors in series is

which is 4 times less than the power delivered across the parallel circuit.

The power is maximum in parallel circuit because the net resistance in the parallel circuit is less than the series circuit and power delivered is inversely proportional to net resistance.